# UBio-MolFM A Universal Molecular Foundation Model for Bio-Systems

### v1

@@ 2.1.2 Relative Energy and Force
UBio-MolFM (S2) 87.975 42.843 2.184
UBio-MolFM (S3) 8.680 16.766 0.832
@@
UBio-MolFM (S2) 142.485 41.325 8.622
UBio-MolFM (S3) 140.162 32.555 13.760
@@
UBio-MolFM (S2) 376.825 36.339 7.343
UBio-MolFM (S3) 336.084 21.637 6.095
@@
UBio-MolFM (S2) 49.920 47.049 30.378
UBio-MolFM (S3) 31.194 20.622 23.552
@@
The transition from Stage 2 to Stage 3 highlights the impact of UBio-Mol26. By adding macromolecular configurations, UBio-MolFM (S3) significantly improves protein-system accuracy, notably reducing force errors in both optimization and MD settings. At the same time, DNA optimization exhibits a regression in $\Delta E$ (8.622 for S2 to 13.760 for S3). This divergence likely reflects imbalanced coverage in UBio-Mol26 (see §4.1), where protein data dominates and nucleic-acid data remain limited to bottom-up fragments. The result underscores a concrete direction for improvement: expand top-down sampling and structural diversity for nucleic acids.
@@ 2.2.1 Solvation Structure
We performed 200 ps $NVT$ simulations for a pure water box at 300 K with both UBio-MolFM (S3) and UMA-S-1p1. We computed the oxygen-oxygen radial distribution function (O–O RDF) and compared against high-precision experimental data [37, 7]. As shown in Figure 3, both models capture the characteristic hydration shells. The two models are closely matched: UMA is slightly closer in the first-peak position, while UBio-MolFM exhibits slightly sharper peak definition. The remaining differences may be influenced by UBio-MolFM’s leaner short-range stack (two local layers).
@@
Figure 3: Structural Fidelity of Liquid Water. Comparison of oxygen-oxygen radial distribution functions (O–O RDF) derived from (a) UBio-MolFM (S3) and (b) UMA-S-1p1 NVTNVT trajectories against experimental references [37, 7]. The two models are closely matched; UMA is slightly closer in the first-peak position, while UBio-MolFM shows slightly sharper peak definition.
@@ 2.2.3 RNA Dynamics
Figure 6 compares Amber99 RNA OL3, UMA-S-1p1, and UBio-MolFM (S3). Amber99 captures the coordination topology (five waters, one OP) but underestimates Mg–O distances and enforces a rigid Mg–O–P angle around 160. UMA-S-1p1 (only $\sim$40 ps trajectory due to simulation efficiency) places the Mg–Owater peak near 2.04 Å but overestimates Mg–OOP distances, yielding an inflated Mg–P distance and a Mg–O–P angle distribution of 120–130, inconsistent with experimental benchmarks.
UBio-MolFM (S3) provides the most accurate structural description. It effectively reproduces the Mg–Owater and Mg–OOP distribution with peak at 2.04 Å, capturing the stronger binding to phosphate oxygens compared with UMA-S-1p1. It also models Mg–O–P angle distributions spanning 120–160, indicating improved flexibility relative to the Amber baseline without the deviations seen in UMA. All three models capture the octahedral O–Mg–O angles ($\approx 90^{\circ}$ and $170^{\circ}$) and the correct coordination number.
@@
Figure 6: Mg2+ Coordination to RNA Phosphate (1L2X). Comparison of geometric distributions for Mg–O distances and angles against experimental benchmarks [48]. Amber99 captures coordination topology but exhibits rigid angles ($\sim$160), while UMA mischaracterizes Mg–phosphate interactions (overestimated Mg–OOP distance). UBio-MolFM (S3) reproduces stronger binding to phosphate oxygens (peak at 2.01 Å) and maintains flexible angles (130–160).
@@ 3.1 Conclusion
However, we also recognize open challenges. While gaining speed over other MLFFs, UBio-MolFM remains substantially slower than classical force fields, limiting current validation to shorter timescales for medium-sized systems. Furthermore, we observed a regression in DNA energy prediction accuracy from Stage 2 to Stage 3, highlighting the need to balance data diversity for nucleic acids. These results position UBio-MolFM not as a final solution, but as a robust proof-of-concept that high-fidelity, generalizable biological simulation is achievable.
@@ 4.3.2 Atom-Balanced Data Loading Mechanism
Atom-Budgeted Sample Packing. Molecular systems exhibit highly diverse sizes. For example, molecules in the OMol25 dataset range from 50 to 300 atoms, while our in-house SVP and TZVPD datasets predominantly span 300–1500 atoms. Such variability leads to severe load imbalance when batching by the number of molecular graphs. To address this issue, we adopt an atom-count–based batching strategy. Specifically, we define a maximum atom budget bs_atom per batch. After randomly shuffling the training set, molecules are greedily added to the current batch until the accumulated atom count reaches bs_atom. Remaining samples are assigned to subsequent batches. This strategy ensures that each GPU processes a similar number of atoms, resulting in more balanced computational cost and improved training efficiency.
@@ 5.2 Code and Inference
https://github.com/IQuestLab/UBio-MolFM/tree/e2formerv2
@@ 5.3 Model Weights
The pretrained weights for the UBio-MolFM series will be released on Hugging Face. We will provide checkpoints from different training stages to support both general chemical modeling and biology-focused downstream applications.
@@ 7  Contributors & Acknowledgments
Acknowledgments
We thank the IQuest Research infrastructure team for providing the computational resources (GPU cluster) required for training and validating UBio-MolFM.

### v2

@@ 2.1.2 Relative Energy and Force
UBio-MolFM (S2) 82.793 43.052 1.923
UBio-MolFM (S3) 8.253 16.511 0.978
@@
UBio-MolFM (S2) 138.955 46.637 8.171
UBio-MolFM (S3) 93.925 44.058 19.454
@@
UBio-MolFM (S2) 460.622 38.379 7.514
UBio-MolFM (S3) 144.730 22.375 5.981
@@
UBio-MolFM (S2) 47.103 46.686 29.579
UBio-MolFM (S3) 40.909 19.783 24.940
@@
The transition from Stage 2 to Stage 3 highlights the impact of UBio-Mol26. By adding macromolecular configurations, UBio-MolFM (S3) dramatically improves protein-system accuracy and RNA optimization, with RNA energy MAE dropping by >60% compared to baselines. Notably, S3 achieves the best force errors and temporal stability ($\Delta E$) in both protein optimization and MD settings. However, DNA optimization remains a challenge, exhibiting a regression in $\Delta E$ (8.171 for S2 to 19.454 for S3) and failing to match the local force precision of baseline models. This divergence likely reflects imbalanced coverage in UBio-Mol26 (see §4.1), where protein and RNA fragments are better represented, while complex DNA configurations remain limited. This result underscores a concrete direction for future research: expanding structural diversity specifically for DNA potential energy surfaces.
@@ 2.2.1 Solvation Structure
We performed 200 ps $NVT$ simulations for a pure water box at 300 K with both UBio-MolFM (S3) and UMA-S-1p1. We computed the oxygen-oxygen radial distribution function (O–O RDF) and compared against high-precision experimental data [37, 7]. As shown in Figure 3, both models capture the characteristic hydration shells with high structural fidelity. UBio-MolFM (S3) demonstrates exceptional agreement with high-precision experimental benchmarks, effectively matching both the first-peak position and the overall structural definition. This performance indicates that UBio-MolFM can recover the delicate hydrogen-bonding network of liquid water at a level competitive with established benchmarks.
@@
Figure 3: Structural Fidelity of Liquid Water. Comparison of oxygen-oxygen radial distribution functions (O–O RDF) derived from (a) UBio-MolFM (S3) and (b) UMA-S-1p1 NVTNVT trajectories against experimental references [37, 7]. Both models exhibit excellent agreement with experiment; UBio-MolFM (S3) accurately reproduces the primary and secondary hydration shells, demonstrating the physical robustness of the learned potential.
@@ 2.2.3 RNA Dynamics
Figure 6 compares Amber99 RNA OL3, UMA-S-1p1, and UBio-MolFM (S3). Amber99 captures the coordination topology (five waters, one OP) but significantly underestimates Mg–O distances ($\sim$1.95 Å) and enforces a rigid Mg–O–P angle around 160, both deviating from high-resolution experimental benchmarks. UMA-S-1p1 (only $\sim$40 ps trajectory due to simulation efficiency) places the Mg–Owater peak near 2.04 Å but overestimates Mg–OOP distances ($\sim$2.20 Å), yielding an inflated Mg–P distance and a Mg–O–P angle distribution centered near 125, inconsistent with experimental coordination geometry.
UBio-MolFM (S3) provides the most accurate structural description. It is the only model that effectively reproduces both the Mg–O distribution peaks at 2.04 Å and the Mg–O–P angle distributions (fluctuating between 120–150) within experimental uncertainty. This suggests that UBio-MolFM accurately captures the site-specific metal binding environment without the over-binding seen in Amber or the structural deviations in UMA. All three models capture the octahedral O–Mg–O angles ($\approx 94^{\circ}$ and $172^{\circ}$) and the correct coordination number.
@@
Figure 6: Mg2+ Coordination to RNA Phosphate (1L2X). Comparison of geometric distributions for Mg–O distances and angles against experimental benchmarks [48]. Amber99 captures coordination topology but exhibits rigid angles ($\sim$160) and underestimates Mg–O distances. UMA mischaracterizes Mg–phosphate interactions with overestimated distances and skewed angles. UBio-MolFM (S3) demonstrates superior fidelity, reproducing both primary Mg–O distances (2.04 Å) and flexible Mg–O–P angles (120–150) within experimental tolerances.
@@ 3.1 Conclusion
However, we also recognize open challenges. While gaining speed over other MLFFs, UBio-MolFM remains substantially slower than classical force fields, limiting current validation to shorter timescales for medium-sized systems. Furthermore, we observed that while absolute energy accuracy improved for DNA structures from Stage 2 to Stage 3, there was a regression in temporal energy stability (local energy differences, $\Delta E$), highlighting the need to further balance data diversity for nucleic acid potential energy surfaces. These results position UBio-MolFM not as a final solution, but as a robust proof-of-concept that high-fidelity, generalizable biological simulation is achievable.
@@ 4.3.2 Atom-Balanced Data Loading Mechanism
Atom-Budgeted Sample Packing. Molecular systems exhibit highly diverse sizes. For example, molecules in the OMol25 dataset range from 50 to 300 atoms, while our in-house SVP and TZVPD datasets span 300–1500 atoms. To rigorously evaluate out-of-distribution (OOD) generalization, we exclude all configurations with more than 1,200 atoms from the training set, reserving the 1,200–1,500 atom regime exclusively for the extrapolation benchmarks described in §2. Such variability leads to severe load imbalance when batching by the number of molecular graphs. To address this issue, we adopt an atom-count–based batching strategy. Specifically, we define a maximum atom budget bs_atom per batch. After randomly shuffling the training set, molecules are greedily added to the current batch until the accumulated atom count reaches bs_atom. Remaining samples are assigned to subsequent batches. This strategy ensures that each GPU processes a similar number of atoms, resulting in more balanced computational cost and improved training efficiency.
@@ 5.2 Code and Inference
https://github.com/IQuestLab/UBio-MolFM
@@ 5.3 Model Weights
The pretrained weights for the UBio-MolFM series are available on Hugging Face at:
https://huggingface.co/IQuestLab/IQuest-UBio-MolFM-V1
We provide checkpoints for Stage 3 to support biology-focused downstream applications.